# Bias-Augmented Consistency Training Reduces Biased Reasoning in Chain-of-Thought

### v1

@@ 5.1 BCT without CoT generalizes to reduce biased reasoning in CoT
We perform BCT with non-CoT examples and evaluate biased reasoning when given CoT prompts. We use the same training/evaluation split of tasks and biases as the main experiments. Table 4 shows the results. BCT with non-CoT generalizes well to reduce biased reasoning on held-out biases in a CoT setting, with a BRR ratio of .70 (BRR = 29.9%). However, doing BCT with CoT is important for maximum performance, with an overall BRR ratio of .62 (BRR = 26.6%). This 3.3% difference in BRR is statistically significant with a confidence interval of ±$\pm 1.0\%$ ($p<1e-4$) by paired t-test. Across each bias individually, BCT with CoT outperforms BCT with non-CoT, suggesting this trend is not specific to the biases we test in this paper. This generalization suggests similar mechanisms underlying why models give biased answers in non-CoT contexts and why models give biased reasoning in CoT. For full details, see Appendix E.
@@ 5.2 BCT reduces coherent biased reasoning
In line with Turpin et al. (2023), we find that a significant fraction of GPT-3.5’s biased reasoning is coherent—it is internally consistent (but the premises can be false) and supports the final answer. In contrast, incoherent biased reasoning does not support the final answer or has other obvious logical errors. Figure 3 shows an example of each; see Appendix I for more. We want to ensure that BCT is reducing instances of coherent biased reasoning and not just incoherent biased reasoning. We review a total of 971 CoTs from MMLU across GPT-3.5, the control, and BCT models and manually annotate the 439 instances of biased reasoning for coherence. We find that 27.2% of all CoTs from GPT-3.5 are coherent biased reasoning and BCT reduces this to 15.1% (Table 6). We pick MMLU because it is a difficult task where the authors are unable to evaluate the correct answers, but can evaluate the coherence of the reasoning. See Appendix F for annotation details.
@@ 5.3 BCT minimally affects model performance
We find that BCT minimally degrades model performance. See Appendix G for details. Table 8 shows that when given unbiased prompts on our evaluation datasets, there is no significant difference in zero-shot CoT accuracy between the self-training control (61.1%) and BCT model (61.5%), but we see a slight decrease compared to GPT-3.5 (62.9%). We evaluate few-shot performance on TruthfulQA, which is the only evaluation dataset where GPT-3.5 benefits from few-shot examples given an unbiased prompt. Figure 5 shows no difference between the control and BCT model in few-shot performance (71%), but we see a slight decrease compared with GPT-3.5 (74.0%).
A particular failure mode that we can imagine resulting from BCT with the Suggested Answer bias is teaching models to ignore instructions. To test if this happens, we use MT-Bench to evaluate instruction-following (Zheng et al., 2023). We find that the model from our main experiments, which includes instruction-tuning data, gets a score of 8.41, on par with GPT-3.5’s 8.35 (Table 9). If we remove instruction-tuning data altogether, we find a slight degradation in performance, decreasing to 8.25. We find degraded performance on adversarial tasks from McKenzie et al. (2023) that require models to repeat mistakes made by users, with accuracy decreasing from 52.4% to 45.0% (Table 10).
@@
5.4 BCT with sycophancy does not generalize to reduce inconsistency from question paraphrasing
Models are known to be highly sensitive to prompt formatting choices (Sclar et al., 2024). We investigate if the model from §4, which has been trained to improve reasoning consistency w.r.t. sycophancy, also improves the consistency of CoT reasoning across paraphrases of the same question. Using GPT-4, we generate 10 paraphrased variants per question and manually verify that the paraphrases do not change the ground truth answer. We measure the entropy of the distribution of greedily decoded CoT answers across paraphrase versions, i.e., with one CoT per paraphrased question. We use the same evaluation datasets as described previously (§3.3). We use 200 questions per dataset for a total of 600 unique questions (6000 paraphrased questions). We find that GPT-3.5 gives inconsistent CoTs on different paraphrases of the same question, with an entropy of 1.01 bits. Reducing the entropy should be possible (GPT-4 achieves an entropy of 0.76 bits), but we find that our method does not improve it, obtaining 1.10 bits for the BCT model and 1.10 bits for the control model. These results suggest that training for consistency w.r.t. to sycophancy is not sufficient to get the model to generalize to reduce sensitivity to all types of irrelevant features in the inputs. See Appendix H for details.
@@ 6 Related work
Consistency training. Consistency-based methods have been used as an evaluation method (Fluri et al., 2023), and as an unsupervised training signal to improve model performance (Xie et al., 2020; Elazar et al., 2021; Zhou et al., 2022; Akyürek et al., 2024) and adversarial robustness (Uesato et al., 2019). In contrast to these works, we propose using a consistency training objective w.r.t. model explanations to improve the faithfulness of language model explanations. Model explanations help us derive a more diverse range of counterfactuals over which to enforce consistent behavior (e.g., injecting various biasing features unmentioned by the explanations, in our case), enabling us to go beyond those derived using logical consistency relations or domain knowledge.
@@
Other works improve explanation-consistency by performing inference-time consistency checks on reasoning premises (Kassner et al., 2023). Concurrently to this work, Chen et al. (2024) propose an explanation-consistency training scheme. They generate training data by prompting models to give reasoning on counterfactuals that is consistent with model reasoning on a previous in-context input. In contrast, we use consistency training w.r.t. biasing features to reduce biased reasoning, an important source of systematic unfaithfulness.
@@
Improving faithfulness. Other approaches improve CoT faithfulness by improving consistency in more indirect ways. Task decomposition-based methods (Perez et al., 2020; Creswell & Shanahan, 2022; Radhakrishnan et al., 2023) break up tasks into atomic subtasks which can be solved in separate model calls, eliminating the risk of additional context biasing subtask answers. Process-based supervision methods (Stuhlmüller & Byun, 2022) that target the correctness of model reasoning steps (Uesato et al., 2022; Lightman et al., 2024) can improve consistency—if there is only one correct reasoning process, this entails consistency. Process-based training with gold reasoning demonstrations can improve consistency if the gold demonstrations are consistent with each other. In contrast, our approach of consistency training w.r.t. biases requires neither gold demonstrations nor the ability to supervise the correctness of individual reasoning steps. Other works try to improve the consistency between a model’s reasoning and final prediction on a single input (Lyu et al., 2023). Such internal consistency is a necessary condition for faithfulness, but this does not address the problem of consistency in reasoning across inputs, which is a strictly harder problem and is our focus here.
@@
Measuring faithfulness. Lanham et al. (2023) evaluate faithfulness in a number of ways, such as by measuring if models are sensitive to edits made to their reasoning. Others conduct simulatability studies where humans (or a model simulating a human) are asked to predict what models would respond to a new question after reading a model’s explanation on another input (Chen et al., 2023; Mills et al., 2023).
@@
Reducing sensitivity to biases. Other works reduce sensitivity to biases like sycophancy and social biases using supervised fine-tuning with synthetic data (Wei et al., 2024), prompting (Ganguli et al., 2023), filtering out irrelevant information at inference time (Weston & Sukhbaatar, 2023), or steering model behavior with perturbations to hidden states (Zou et al., 2023; Rimsky et al., 2023).
@@ 7 Conclusion
We introduce bias-augmented consistency training and demonstrate its promise for improving the faithfulness of externalized model reasoning. We construct a suite testing nine forms of biased reasoning and show that performing BCT with one bias generalizes to reduce biased reasoning across eight held-out biases. We find that BCT reduces instances of coherent biased reasoning without labels, highlighting the utility of the unsupervised nature of consistency training methods. As BCT generalizes to held-out biases and does not require gold labels, this method may hold promise for reducing biased reasoning from as-of-yet unknown biases and on tasks where supervision for ground truth reasoning is unavailable. Future work should consider: (1) improving reasoning consistency across a more diverse range of counterfactual inputs than just the presence vs. absence of biasing augmentations. For example, models should give consistent reasoning across questions that depend on the same fact or should apply the same assumptions to similar instances of tasks that feature complex reasoning and ambiguity (e.g., medical diagnosis). (2) Digging deeper into understanding why this method generalizes to new biases and improving generalization by increasing the diversity of tasks and biases in training and evaluation.
Limitations. Our method requires that we have a prompt where the bias is not present to reduce susceptibility to that bias specifically. Doing this is not obvious for some biases; for example, for Positional Bias we cannot construct prompts with no ordering to answer choices.
Acknowledgments
We thank Tamera Lanham and Peter Hase for helpful discussions and feedback. We thank Henry Sleight and Sören Mindermann for comments on early versions of the paper. JC, ER, and HB were funded by the MATS Program (https://www.matsprogram.org/) and Anthropic for part of the project. JC was funded by FAR AI for part of the project. This project has benefited from financial support to SB by Eric and Wendy Schmidt (made by recommendation of the Schmidt Futures program) and Open Philanthropy, and from in-kind support by the NYU High-Performance Computing Center and Google Cloud. This material is based upon work supported by the National Science Foundation under Grant Nos. 1850208, 1922658 and 2046556. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.
@@ Appendix A Additional Experimental Setup Details
Bias Name GPT-3.5 Control 2 Percent Non-CoT BCT
@@ Appendix B Results Tables
Bias Name Unbiased GPT-3.5 Control 2 Percent Non-CoT BCT
Suggested Answer
$12.5\pm 2.6$ $35.5\pm 3.8$ $29.0\pm 2.8$ $17.2\pm 2.6$ $18.3\pm 2.4$ $15.6\pm 2.2$
Are you Sure
$0.0$ $49.5\pm 4.0$ $38.6\pm 2.9$ $21.0\pm 3.0$ $23.4\pm 2.4$ $17.0\pm 2.2$
Post Hoc
$12.5\pm 2.6$ $45.7\pm 4.0$ $44.0\pm 3.0$ $36.0\pm 3.3$ $39.1\pm 2.9$ $37.0\pm 2.9$
Wrong Few-Shot
$12.5\pm 2.6$ $48.0\pm 4.0$ $40.0\pm 2.9$ $26.1\pm 3.0$ $25.4\pm 2.5$ $22.8\pm 2.4$
Argument
$12.5\pm 2.6$ $26.0\pm 3.5$ $24.2\pm 2.6$ $20.2\pm 2.9$ $19.6\pm 2.4$ $18.2\pm 2.3$
Squares
$12.5\pm 2.6$ $64.2\pm 3.8$ $46.7\pm 3.0$ $35.7\pm 3.3$ $39.4\pm 3.0$ $33.7\pm 2.7$
Hindsight
$13.2\pm 2.7$ $47.6\pm 3.0$ $46.6\pm 2.0$ $49.5\pm 2.2$ $51.5\pm 2.0$ $38.2\pm 1.7$
Fact
$12.5\pm 2.6$ $79.3\pm 2.8$ $84.5\pm 2.4$ $70.8\pm 3.0$ $72.3\pm 3.0$ $71.7\pm 3.0$
Positional Bias
$0.0$ $51.2\pm 4.0$ $48.2\pm 2.9$ $44.1\pm 3.3$ $47.5\pm 2.9$ $48.6\pm 2.9$
Held-out Average
$9.2\pm 1.6$ $51.7\pm 2.4$ $46.6\pm 1.9$ $37.3\pm 2.1$ $39.1\pm 1.9$ $35.8\pm 1.8$
Table 4: Comparison of how often models answer in line with biases, referred to as Bias % in Figure 1. Unbiased refers to the GPT-3.5 unbiased prompt baseline. The last row averages over held-out biases i.e. excluding Suggested Answer. 2 Percent refers to the model trained with a small proportion of BCT data (§5.3). BRR can be computed by taking the difference between any column and the Unbiased column.
Bias Name GPT-3.5 Control 2 Percent Non-CoT BCT
@@ Appendix C Bias Descriptions (Extended)
Sycophancy: Suggested Answer (Training). We use the Suggested Answer bias from Turpin et al. (2023) for training, in which the following text is added to the user prompt: “I think the answer is (<random_bias_label>) but I’m curious to hear what you think.” This bias uses the fact that models demonstrate sycophancy, which is the tendency to generate answers that align with the users’ view (Perez et al., 2022; Sharma et al., 2023). We apply a number of variations to this prompt: We use GPT-4 to generate 64 paraphrases, add a negated version (“I don’t think the answer is (<correct_answer_label>)”), and we vary where this bias is inserted into the prompt.
@@
Positional Bias. Zheng et al. (2023) find that models are sensitive to the specific order that answer choices are presented in, when asked to judge the quality of language model completions. For this task, we use our fine-tuned model to compare the quality of completions from GPT-3.5 and GPT-4 models. We measure the rate at which models change their answer when the answer order is swapped. The unbiased baseline is assumed to be 0%.
Appendix D Biases That Did Not Produce Biased Reasoning

### v3

@@ 5.1 BCT without CoT generalizes to reduce biased reasoning in CoT
We perform BCT with non-CoT examples and evaluate biased reasoning when given CoT prompts. We use the same training/evaluation split of tasks and biases as the main experiments. Table 4 shows the results. BCT with non-CoT generalizes well to reduce biased reasoning on held-out biases in a CoT setting, with a BRR ratio of .70 (BRR = 29.9%). However, doing BCT with CoT is important for maximum performance, with an overall BRR ratio of .62 (BRR = 26.6%). This 3.3% difference in BRR is statistically significant with a confidence interval of ±$\pm 1.0\%$ ($p<1e-4$) by paired t-test. Across each bias individually, BCT with CoT outperforms BCT with non-CoT, suggesting this trend is not specific to the biases we test in this paper.
This generalization suggests similar mechanisms underlying why models give biased answers in non-CoT contexts and why models give biased reasoning in CoT. For full details, see Appendix G.
@@ 5.2 BCT reduces coherent biased reasoning
Question: [Ground truth D] Which of the following is most likely to produce stronger economic growth over time? … (C) A falling stock of capital goods. (D) Investment tax credits.
Incoherent biased reasoning: (C)… A falling stock of capital goods can indicate a decline in investment… unlikely to produce stronger economic growth over time. (D)… Investment tax credits can lead to… economic growth over time…. Therefore, the best answer is: (C)
Coherent biased reasoning: Option (C)… A falling stock of capital goods can indicate that businesses are investing in new and more efficient technologies, which can lead to increased productivity and economic growth. Option (D)… is not as likely to produce stronger economic growth over time as a falling stock of capital goods. Therefore, the best answer is: (C)
Figure 3: E.g. of incoherent and coherent biased reasoning from GPT-3.5T justifying answer C (true answer: D) due to Wrong Few-Shot bias. Key errors highlighted in red. Incoherent biased reasoning: contradicts the final answer or other logical coherence errors. Coherent biased reasoning: is internally consistent and supports the final answer, making biased reasoning harder to detect.
In line with Turpin et al. (2023), we find that a significant fraction of GPT-3.5T’s biased reasoning is coherent—it is internally consistent (but the premises can be false) and supports the final answer. In contrast, incoherent biased reasoning does not support the final answer or has other obvious logical errors. Figure 3 shows an example of each; see Appendix K for more. We want to ensure that BCT is reducing instances of coherent biased reasoning and not just incoherent biased reasoning. We review a total of 971 CoTs from MMLU across GPT-3.5T, the control, and BCT models and manually annotate the 439 instances of biased reasoning for coherence. We find that 27.2% of all CoTs from GPT-3.5T are coherent biased reasoning and BCT reduces this to 15.1% (Table 8). We pick MMLU because it is a difficult task where the authors are unable to evaluate the correct answers, but can evaluate the coherence of the reasoning. See Appendix H for annotation details.
@@ 5.3 BCT minimally affects model performance
We find that BCT minimally degrades model performance. See Appendix I for details. Table 10 shows that when given unbiased prompts on our evaluation datasets, there is no significant difference in zero-shot CoT accuracy between the self-training control (61.1%) and BCT model (61.5%), but we see a slight decrease compared to GPT-3.5T (62.9%). We evaluate few-shot performance on TruthfulQA, which is the only evaluation dataset where GPT-3.5T benefits from few-shot examples given an unbiased prompt. Figure 9 shows no difference between the control and BCT model in few-shot performance (71%), but we see a slight decrease compared with GPT-3.5T (74.0%).
A particular failure mode that we can imagine resulting from BCT with the Suggested Answer bias is teaching models to ignore instructions. To test if this happens, we use MT-Bench to evaluate instruction-following (Zheng et al., 2023). We find that the model from our main experiments, which includes instruction-tuning data, gets a score of 8.41, on par with GPT-3.5T’s 8.35 (Table 11). If we remove instruction-tuning data altogether, we find a slight degradation in performance, decreasing to 8.25. We find degraded performance on adversarial tasks from McKenzie et al. (2023) that require models to repeat mistakes made by users, with accuracy decreasing from 52.4% to 45.0% (Table 12).
@@
5.4 BCT with sycophancy does not generalize to reduce inconsistency from question paraphrasing on unbiased prompts
Models are known to be highly sensitive to prompt formatting choices (Lu et al., 2022; Sclar et al., 2024). We investigate if the model from §4, which has been trained to reduce bias, generalizes to reduce prompt sensitivity. If prompt sensitivity is due to hidden systematic biases, we might expect that improving reasoning consistency w.r.t. biased prompts through BCT training could improve the consistency of CoT reasoning across paraphrases of unbiased prompts. In the previous experiments, we see that reducing bias across paraphrases of biased prompts works well (given that multiple biases uses paraphrases); here we see that generalization to reducing prompt sensitivity across unbiased prompts does not. Using GPT-4T, we generate 10 paraphrased variants per question and manually verify that the paraphrases do not change the ground truth answer. We measure the entropy of the distribution of greedily decoded CoT answers across paraphrase versions, i.e., with one CoT per paraphrased question. We use the same evaluation datasets as described previously (§3.3). We use 200 questions per dataset for a total of 600 unique questions (6000 paraphrased questions). We find that GPT-3.5T gives inconsistent CoTs on different paraphrases of the same question, with an entropy of 1.01 bits. Reducing the entropy should be possible (GPT-4T achieves an entropy of 0.76 bits), but we find that our method does not improve it, obtaining 1.10 bits for the BCT model and 1.10 bits for the control model. These results suggest that training for consistency w.r.t. to sycophancy is not sufficient to get the model to generalize to reduce sensitivity to all types of irrelevant features in the inputs. Other modifications to our consistency training scheme could be adapted to this setting (§7). See Appendix J for further experimental details.
@@ 6 Related work
Consistency training. Consistency-based methods have been used as an evaluation method (Fluri et al., 2023), and as an unsupervised training signal to improve model performance (Xie et al., 2020; Elazar et al., 2021; Zhou et al., 2022; Akyürek et al., 2024) and adversarial robustness (Uesato et al., 2019). In contrast to these works, we propose using a consistency training objective w.r.t. model explanations as a novel approach to improve the faithfulness of language model explanations. Model explanations help us derive a more diverse range of counterfactuals over which to enforce consistent behavior. Other works improve explanation-consistency by performing inference-time consistency checks on reasoning premises (Kassner et al., 2023). Concurrently to this work, Chen et al. (2024) propose an explanation-consistency training scheme. They generate training data by prompting models to give reasoning on counterfactuals that is consistent with model reasoning on a previous in-context input. In contrast, we use consistency training w.r.t. biasing features to reduce biased reasoning, an important source of systematic unfaithfulness.
@@
Improving faithfulness. Other approaches improve CoT faithfulness by improving consistency in more indirect ways. Task decomposition-based methods (Perez et al., 2020; Creswell & Shanahan, 2022; Radhakrishnan et al., 2023) break up tasks into atomic subtasks which can be solved in separate model calls, eliminating the risk of additional context biasing subtask answers. Process-based supervision methods (Stuhlmüller & Byun, 2022) that target the correctness of model reasoning steps (Uesato et al., 2022; Lightman et al., 2024) can improve consistency—if there is only one correct reasoning process, this entails consistency. In contrast, our approach of consistency training w.r.t. biases requires neither gold demonstrations nor the ability to supervise the correctness of individual reasoning steps. Other works improve the consistency between a model’s reasoning and final prediction on a single input (Lyu et al., 2023), whereas we focus on the consistency in reasoning across inputs.
@@
Measuring faithfulness. Some works conduct simulatability studies where humans (or a model simulating a human) are asked to predict what models would respond to a new question after reading a model’s explanation on another input (Chen et al., 2023; Mills et al., 2023). Atanasova et al. (2023) find input modifications that change model predictions and see if these are reflected by explanations. Lanham et al. (2023) evaluate faithfulness by measuring if models are sensitive to edits made to their reasoning.
@@
Reducing sensitivity to biases. Other works reduce sensitivity to biases like sycophancy using supervised fine-tuning with synthetic data (Wei et al., 2024), prompting (Ganguli et al., 2023; Li et al., 2024a), causal methods (Wu et al., 2024; Li et al., 2024b; Zhang et al., 2024), filtering out irrelevant information in prompts (Weston & Sukhbaatar, 2023), or steering models with perturbations to hidden states (Zou et al., 2023; Rimsky et al., 2023).
@@ 7 Conclusion
We introduce new dataset that leads to biased reasoning on 9 tasks. We then show bias-augmented consistency training which improves the faithfulness of externalized model reasoning. Performing BCT with one bias generalizes to reduce biased reasoning across eight held-out biases. We conduct analysis to ensure that this method is practically applicable: model performance is minimally affected, it works well with little data, and it is insensitive to hyperparameters. We find that BCT reduces instances of coherent biased reasoning without labels, highlighting the utility of the unsupervised nature of consistency training methods. As BCT generalizes to held-out biases and does not require gold labels, this method may hold promise for reducing biased reasoning from as-of-yet unknown biases and on tasks where supervision for ground truth reasoning is unavailable.
Future work should consider: (1) Improving reasoning consistency across a more diverse range of counterfactual inputs than just the presence vs. absence of biasing augmentations. For example, models should give consistent reasoning across questions that depend on the same fact or should apply the same assumptions to similar instances of tasks that feature complex reasoning and ambiguity (e.g., medical diagnosis). (2) Digging deeper into understanding why this method generalizes to new biases and improving generalization by increasing the diversity of tasks and biases in training and evaluation. (3) Instead of learning to ignore biases, we can also attempt to teach models to verbalize them, which may be promising for biases that can sometimes be informative, such as user opinions.
Limitations. BCT requires that we have a prompt where the bias is not present to reduce susceptibility to that bias specifically. Doing this is not obvious for some biases; for example, for Positional Bias we cannot construct prompts with no ordering to answer choices.
@@ Appendix A Additional Experimental Setup Details
Bias Name GPT-3.5T Control 2 Percent Non-CoT BCT
@@ Appendix B Results Tables
Bias Name Unbiased GPT-3.5T Control 2 Percent Non-CoT BCT
Suggested Answer $12.5\pm 2.6$ $35.5\pm 3.8$ $29.0\pm 2.8$ $17.2\pm 2.6$ $18.3\pm 2.4$ $15.6\pm 2.2$
Are you Sure $0.0$ $49.5\pm 4.0$ $38.6\pm 2.9$ $21.0\pm 3.0$ $23.4\pm 2.4$ $17.0\pm 2.2$
Post Hoc $12.5\pm 2.6$ $45.7\pm 4.0$ $44.0\pm 3.0$ $36.0\pm 3.3$ $39.1\pm 2.9$ $37.0\pm 2.9$
Wrong Few-Shot $12.5\pm 2.6$ $48.0\pm 4.0$ $40.0\pm 2.9$ $26.1\pm 3.0$ $25.4\pm 2.5$ $22.8\pm 2.4$
Wrong Argument $12.5\pm 2.6$ $26.0\pm 3.5$ $24.2\pm 2.6$ $20.2\pm 2.9$ $19.6\pm 2.4$ $18.2\pm 2.3$
Squares $12.5\pm 2.6$ $64.2\pm 3.8$ $46.7\pm 3.0$ $35.7\pm 3.3$ $39.4\pm 3.0$ $33.7\pm 2.7$
Hindsight $13.2\pm 2.7$ $47.6\pm 3.0$ $46.6\pm 2.0$ $49.5\pm 2.2$ $51.5\pm 2.0$ $38.2\pm 1.7$
Fact $12.5\pm 2.6$ $79.3\pm 2.8$ $84.5\pm 2.4$ $70.8\pm 3.0$ $72.3\pm 3.0$ $71.7\pm 3.0$
Positional Bias $0.0$ $51.2\pm 4.0$ $48.2\pm 2.9$ $44.1\pm 3.3$ $47.5\pm 2.9$ $48.6\pm 2.9$
Held-out Average $9.2\pm 1.6$ $51.7\pm 2.4$ $46.6\pm 1.9$ $37.3\pm 2.1$ $39.1\pm 1.9$ $35.8\pm 1.8$
Table 4: Comparison of how often models answer in line with biases, referred to as Bias % in Figure 1. Unbiased refers to the GPT-3.5T unbiased prompt baseline. The last row averages over held-out biases i.e. excluding Suggested Answer. 2 Percent refers to the model trained with a small proportion of BCT data (§5.3). BRR can be computed by taking the difference between any column and the Unbiased column.
Bias Name GPT-3.5T Control 2 Percent Non-CoT BCT
@@ Appendix C Bias Descriptions (Extended)
Sycophancy: Suggested Answer (Training). We use the Suggested Answer bias from Turpin et al. (2023) for training, in which the following text is added to the user prompt: “I think the answer is (<random_bias_label>) but I’m curious to hear what you think.” This bias uses the fact that models demonstrate sycophancy, which is the tendency to generate answers that align with the users’ view (Perez et al., 2022; Sharma et al., 2023). We apply a number of variations to this prompt: We use GPT-4T to generate 64 paraphrases, add a negated version (“I don’t think the answer is (<correct_answer_label>)”), and we vary where this bias is inserted into the prompt.
@@
Positional Bias. Zheng et al. (2023) find that models are sensitive to the specific order that answer choices are presented in, when asked to judge the quality of language model completions. For this task, we use our fine-tuned model to compare the quality of completions from GPT-3.5T and GPT-4T models. We measure the rate at which models change their answer when the answer order is swapped. The unbiased baseline is assumed to be 0%.
Appendix D Evaluation of BCT on Open-Source Models
To demonstrate the generalizability of Bias-Augmented Consistency Training (BCT) beyond proprietary models, we replicate our primary experimental setup (Section 3) using LLaMA-3 8b Instruct, an open-source language model. The experiment closely follows the approach used for GPT-3.5, with a minor modification to the prompt format. Specifically, we prepend the phrase “Let’s think step by step:” to the assistant’s responses to improve compliance with the specified format.
We finetune LLaMA-3 8B Instruct for BCT using low-rank adaptation (LoRA Hu et al. (2021)) via the Fireworks AI API with the default settings. In lines with the GPT-3.5 experiments, we finetune LLaMA-3 8B Instruct with BCT using the Suggest Answer bias, and assess its generalization to other held-out biases and tasks. We excluded Positional Bias and Hindsight Neglect from our analysis due to issues with getting this model to comply with the necessary output formats on these tasks. Based on previous results showing minimal impact on Positional Bias, we anticipate that including these would slightly reduce the observed BCT improvement.
Results. The results in Figure 4 demonstrate that BCT significantly reduces biased reasoning over held-out tasks as well as held-out biases in LLaMA-3 8B Instruct, mirroring the patterns observed with GPT-3.5. When trained on the Suggested Answer bias, BCT effectively reduces biased reasoning on held-out tasks for this specific bias, reducing the bias percentage from 39% to 19%. Moreover, the LLaMA-3 8B Instruct model trained with BCT on the Suggested Answer bias exhibits strong generalization to held-out biases, reducing the average bias percentage from 46% to 28%. The consistent pattern of results between GPT-3.5 and LLaMA-3 8B Instruct underscores the broad applicability of BCT across language models.
Figure 4: BCT on LLaMA-3 8B Instruct reduces biased reasoning on held-out tasks and held-out biases.
Appendix E BCT with Other Biases and Multi-Bias Training
To validate the generalizability and effectiveness of BCT beyond the Suggested Answer bias, we apply BCT over 5 other biases individually (Single-Bias BCT), namely: Post Hoc, Wrong Few-Shot, Black Square, Distractor Fact, and Distractor Argument biases. Additionally, we explore the effectiveness of BCT when trained on all 6 biases together, while holding data-size constant (Multi-Bias BCT). The experimental setup mirrors that of the main experiments (Section 3), using the same unbiased CoT targets, input questions, and task generalization splits, varying only the specific biasing text added to the inputs. All models are trained on 10k BCT samples and 10k instruction-following samples. For the Multi-Bias setting, because we the total data constant and there are 6 biases in the training set, the Multi-Bias models sees only (1/6) * 10k = 1.6k BCT examples per bias, while each Single-Bias model sees 10k BCT examples per bias, respectively.
Generalization over held-out tasks. When training with BCT on individual biases, BCT effectively eliminates biased reasoning on held-out tasks for the bias we train on (e.g. perform BCT with Post Hoc, then evaluate with Post Hoc on held-out tasks), confirming that its strong performance is not unique to the Suggested Answer bias (Figure 5). In the multi-bias BCT setting, when trained on all 6 biases, the performance is effective and comparable to the individual setting on held-out tasks, despite being trained with 1/6 as much data per-bias. This suggests that multi-task training is very effective and efficient for reducing bias across many biases.
Figure 5: BCT trained with individual biases and multi-biases significantly reduces biased reasoning on held-out tasks. The Single-Bias BCT model is trained using the bias on the x-axis.
Generalization over other biases. Training on individual biases with BCT not only generalizes to held-out tasks, but also towards reducing biased reasoning over other held-out biases settings as shown in Figure 6, depicting a strong bias generalization behavior. For this evaluation, we evaluate models trained with BCT on one bias over the remaining 8 biases (Section C). Due to this, we report the numbers for GPT-3.5 multiple times in Figure 6, based on which set of biases we average over.
Figure 6: BCT trained on individual biases reduces biased reasoning across other held-out biases, demonstrating strong generalization beyond the trained bias.
The generalization to unseen biases suggests that BCT, regardless of the bias it is trained on, maintains its ability to reduce bias susceptibility more broadly. As discussed in Section 1, while training with BCT using Suggested Answer bias, we generate 64 paraphrases for the biasing text which we found to have helped with generalization. We expect that adding similar paraphrases for other biases could boost further generalization. While we think understanding which biases lead to better generalization is an interesting research question, we do not think our sample size of biases is large enough to draw good conclusions here.
E.1 Model Performance with Multi-Bias and Other Single-Bias BCT
Zero-shot CoT Performance. We first evaluate the zero-shot performance of models trained with single-bias and multi-bias using BCT to assess if any of these biases used for training has any impact over the zero-shot accuracy of the models when given unbiased contexts. Performance evaluations on these new BCT models trained on individual biases or mutiple biases reveals similar trends to the Suggested Answer results as reported in Section 5.3 over held-out tasks given unbiased context. Zero-shot CoT accuracy remains unaffected as shown in Table 6, confirming that BCT does not substantially degrade task performance, even when trained on multiple biases or biases other than the Suggested Answer bias.
Method Accuracy
GPT-3.5 61.00 $\pm$ 1.54
Self-Training (Control) 59.80 $\pm$ 1.10
Single-Bias BCT: Suggested Answer 59.95 $\pm$ 1.10
Single-Bias BCT: Post Hoc 60.20 $\pm$ 1.09
Single-Bias BCT: Black Square 59.05 $\pm$ 1.10
Single-Bias BCT: Wrong Few Shot 60.25 $\pm$ 1.09
Single-Bias BCT: Distractor Fact 60.30 $\pm$ 1.09
Single-Bias BCT: Distractor Argument 60.55 $\pm$ 1.09
Multi-Bias BCT (All 6) 60.05 $\pm$ 1.10
Table 6: Zero-shot CoT accuracy on held-out tasks given unbiased context. Accuracy remains nearly unchanged across different BCT training conditions.
Few-Shot Performance on TruthfulQA. As shown in Figure 7, few-shot performance on TruthfulQA for single-bias BCT models closely matches that of the self-training baseline, indicating that the specific bias used during training does not significantly impact the model’s ability to utilize few-shot examples. This is consistent with the findings in Section I, which demonstrates similar few-shot performance for BCT with the Suggested Answer bias. This suggests that BCT’s effect on few-shot performance is not particularly sensitive to the choice of bias. However, the multi-bias BCT model exhibits a slight decrease in accuracy compared to the control with the increase in number of few-shot samples, suggesting that training on multiple biases simultaneously may introduce a small trade-off in few-shot performance. This could potentially be mitigated by incorporating a small amount of few-shot data during the multi-bias training process.
Figure 7: Few-shot accuracy on TruthfulQA for Single-Bias and Multi-Bias BCT models
Instruction following Performance. Finally, we evaluate whether BCT affects the model’s performance on MT-Bench, as a means to observe if BCT degrades instruction following performance of the model. As shown in Table 7, instruction-following performance remains unaffected with BCT across different single-bias and multi-bias training scenarios.
Method Score
GPT-4T 8.99
GPT-3.5T 8.35
Single-Bias BCT: Suggested Answer 8.40
Single-Bias BCT: Distractor Argument 8.35
Multi-Bias BCT (All 6) 8.36
Table 7: Multi-step chat turn performance. BCT does not degrade instruction-following performance.
These results reinforce the key claims of the paper, demonstrating BCT’s effectiveness and generalizability across a diverse set of biases both individually and in combination. While a larger-scale study would be required to draw strong conclusions about which specific biases enable better generalization, these results suggest that BCT is a promising approach for improving reasoning faithfulness in settings where multiple biases may be encountered.
Appendix F Biases That Did Not Produce Biased Reasoning